# Theorem-Carrying-Transaction: Runtime Certification to Ensure Safety for Smart Contract Transactions

### v1

@@ 7.2 Complexity of the hypotheses and math
tokenA.totalSupply<$\texttt{2}^{\texttt{255}}$& tokenB.totalSupply<$\texttt{2}^{\texttt{255}}$
@@
tokenB. totalSupply<$\texttt{2}^{\texttt{255}}$& tokenA.totalSupply<$\texttt{2}^{\texttt{255}}$
@@
Each of these hypotheses contains the condition pair.token0 == tokenB && pair.token1 == tokenA. Actually, there are three dual-hypotheses, which Table II does not show due to the space constraint. They contain the condition pair.token0 == tokenA && pair.token1 == tokenB. Hence, each pair contract needs six theorems to cover normal liquidity and swap transactions.
The hypotheses and math reveal the precise relation between the design-level properties and the code. There are two discrepancies to bridge. First, the formulas in Table II contain many divisions, but the code performs integer arithmetic. Concretely checking conditions $x_{2}/x_{1}=y_{2}/y_{1}$ and $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$ in Figure 9 will result in false due to the accumulated imprecision of integer operations. Second, for every swap transaction, a small processing fee is charged by the pair as the incentive for the LPs. The rate is set by factory.swapFeeRate, which is usually set to 3 to specify a 0.3% rate. The Swap code trace is equivalent to the formula in Table II only when the fee rate equals 0 precisely. Therefore, the common wisdom that “Uniswap’s swap transaction obeys the property $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$” is not precise. The actual meaning is the conjunction of two statements: (1) the code trace of the transaction is identical to the code trace of a zero-fee-rate transaction, (2) the latter trace, when assuming the real number arithmetic, implies the precise equality $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$. These statements can only be checked symbolically. In TCT, the path hash mechanism and Boogie’s data type real are used. They bridge the non-trivial gap between the design and the code.
7.3 Runtime overhead
The measurements in Table III and Table IV are made using an HP Z4 G4 Workstation with a 3.80 GHz CPU and 32 GB RAM. Table III shows the runtime overhead on Geth 1.13. The Rntrcy transaction is not included because we intentionally make a false theorem for this case. Also, we do not include the five transfer transactions in the measurements because they are very similar to TrfrProxy. We use the Golang utility time.Now().UnixNano() for the measurement. The baseline is the unmodified Geth. Every measurement is repeated using 20 transactions. Each cell reports the average and the percentage of the standard deviation over the average.
The TCT overhead comes from the hypothesis checking and the path hash computation. For each transaction, these two computations are repeated $10^{6}$ times for higher precision of the measurement. The overhead is the sum of these two numbers divided by the baseline. TrfrProxy and NoRntrcy are simple token transactions. The overheads are both 0.20%. As a comparison, Solythesis [13] is a Solidity code instrumentation technique to insert runtime checks to enforce token properties similar to Prop. 1. It reports 24% slowdown in transaction processing.
TABLE III: Overhead on Geth 1.13 (in nanoseconds).
@@ 7.3 Runtime overhead
TABLE IV: Concolic execution time and proof time.
@@
Table IV reports the concolic execution time and the Boogie proof time. Compared to typical formal verification tasks, these numbers are quite low because the reasoning is only about a straight-line code. On the other hand, they are five orders of magnitude higher than those in Table III. The theorem repo turns them into a one-time cost rather than a per-transaction cost. For normal transactions that the programmers anticipate, this one-time cost is paid at the testing time before the contracts process real user transactions.
7.4 Number of theorems expected
The transactions entering from transferProxy of MultiVulnToken and from transfer of BNB, USDC, stETH and TONCOIN require only one theorem, as there is no branch. Function transfer of USDT has a branch, so it needs two theorems (actually, one theorem with two path hashes). Each Uniswap transaction entering from Swap, AddLiquidity, or RemoveLiquidity needs two dual theorems for each token-pair, as explained earlier. In other words, the total number of theorems needed is six times the total number of the pair contracts in order to cover these three operations. Having a linear relation is important for TCT to be applicable to a given scenario. We show that this is true for the studied cases. In Section 8, we propose future work for maintaining the linear relation between the number of theorems and the number of contracts for more complex potential scenarios.
7.5 Insightful violation conditions
@@
TCT’s process explicates insightful conditions, which are often below the design level, such as tx.origin != pair and to != pair for Swap. The former states that the swap transaction must not be issued by the pair itself, and the latter states that the account receiving the swapped-out token from the pair must not be the pair itself. When the conditions are violated, the inverse proportionality will not hold. To show this, we issue six swap transactions to Geth. The reserves of the pair (i.e., $x$ and $y$) are shown in Table V. The former three transactions conform to to != pair, and the latter three violate it. The deviation is the current $x\times y$ divided by the previous $x\times y$ then subtracting 1. Inverse proportionality implies zero deviation. Because of the 0.3% fee, the deviations are [0.0008 , 0.0010] in the second and third rows. However, with to = pair for the last three rows, the value of $y$ becomes unchanged, because the amount of the swapped-out token goes back to the pair. This makes the deviation very large (in fact, arbitrarily large). Hence, if an investment strategy depends on Uniswap unconditionally ensuring inverse proportionality, it can be dangerous. This is similar to integer overflow and reentrancy, as the condition is a detail not foreseeable at the design level. With TCT, specifying inverse proportionality by the designer leads to the explication of conditions like this.
TABLE V: Inverse proportionality violated when to=pair.
$x$ $y$ $x\times y$ Deviation
2000 1000 2000000 N/A
2010 996 2001960 0.0010
2030 987 2003610 0.0008
2050 987 2023350 0.0099
2070 987 2043090 0.0098
2190 987 2161530 0.0580
8 Roadmap Ahead
TCT is a long-term vision for the community. Our prototype shows that it points to a promising direction. Two sub-problems need significant emphasis when we deploy TCT in the real-world setting.
8.1 Consensus of theorem-proving
@@
The Ethereum nodes need to have a consensus on which theorems have been proven. Boogie’s underlying SMT prover is Z3 [31], which is deterministic (unless timeout-based proving tactics are enabled). Similar to Ethereum’s gas mechanism for transaction execution, a prover-gas mechanism should be designed so that all Z3 provers will yield the same result: “sat”, “unsat” or “out of prover gas” for a given SMT problem. Note that some slow nodes falling behind the pace of the majority is a normal situation in Ethereum’s operation, which does not prevent the consensus to be formed. Another potential mechanism to consider is to create a dedicated prover-node consortium, which forms the consensus about which theorems have been proven. Normal Ethereum nodes no longer need to run Z3. Their theorem repositories only need to sync with the latest published snapshot of the consortium’s repository.
8.2 Carrying multiple theorems
In the scenarios that we have discussed, the number of theorems and the number of contracts have a nice linear relation. However, we can imagine scenarios with an exponential number of paths corresponding to independent program branches. We introduce two concepts: “summary” and “carrying multiple theorems”.
Branches often cause local divergence. These so-called diamond splitting can be eliminated by “summaries”. For example, the splitting “if (b) a=1; else a=2;” can be summarized as “modifies(a); assume(a==1 || a==2);”, meaning that no variable other than a is modified, and a is set to either 1 or 2. Reference [32] introduces the notion of summaries for symbolic execution.
Currently, every transaction only carries one theorem. It can be extended to multiple theorems for a transaction shown in Figure 10. ContractA::FuncA calls three cross-contract functions sequentially. The transaction can carry theorem2, theorem3 and theorem4 to prove the postconditions of the three functions. Then, these postconditions are used as the summaries of the functions when proving theorem1 about the safety of running ContractA::FuncA. With this mechanism, the whole transaction is broken into four traces, so the combinatorial complexity is avoided.
Figure 10: A transaction to be proven by multiple theorems.
9 Related Work
TCT is inspired by PCC, PCA and concolic testing. In addition, we summarize the following related technologies.
Static verification for smart contracts. Oyente [33] and Mythril [34] are symbolic verifiers for EVM bytecode. Solidity SMT Checker [35] is a feature built into the Solidity compiler to check assertions and alert possibilities of reentrant calls. Halmos [36] is a recent symbolic analyzer for EVM bytecode compiled from Solidity. In our prototype, we consulted the Halmos code for the EVM semantics.
@@ 9 Related Work
VerX [37] is a static verifier to prove temporal properties for a set of contracts, with certain assumptions about the “external” contracts that they interact with. The temporal logic is focused on properties that correlate contract states across different transactions. It can reason about liveness properties, which TCT is unable to. Annotary [38] is a Mythril-based verifier that can examine cross-contract transactions. To tackle the state explosion problem, Annotary uses different strategies to treat different variables as either concrete or symbolic, depending on their data types. Note that the authors refer to this approach as “concolic execution”. The meaning is very different from ours.
As explained in the introduction, these static verification techniques consider one or a few “contracts of interest” as the program. Without the runtime info of the contract reference topology, any function making a cross-contract call has arbitrary semantics because the reference is essentially dangling, so some assumptions have to be made about the unknown. Moreover, modeling hypothetical malicious users is a job on thin ice, which leads to false positives and false negatives. These techniques provide offline advisories to programmers, unlike TCT’s runtime certification goal.
@@
Code instrumentation. The Solythesis approach [13] is mentioned earlier. It instruments Solidity code to insert “delta updates” so that aggregation properties, such as sum-of-map, count-of-subset, can be calculated with much lower cost than naively computing the aggregations. The idea is to add auxiliary arrays and maps to keep track of incremental updates of the aggregation quantities when the original code updates the corresponding array/map elements. It is unclear whether Solythesis handles multi-contract situations, as the evaluation is only conducted on 23 individual token contracts. The instrumented contracts show a 24% slowdown.
Other related techniques. The essence of concolic approaches is to take advantage of the logic equivalence between different concrete executions. Some techniques share this essence. Forerunner [39] utilizes the logic equivalence of code traces and data dependencies to enable Ethereum to perform speculative execution for transaction processing. CST [40] is a technique to prove global properties for multi-party web protocol programs, such as OAuth and single-sign-on code. Every a protocol execution reassembles the sequence of the function-calls in the involved parties. The sequence is symbolically examined against the multi-party properties to prove its security or expose bugs.
10 Conclusion
Trustworthiness of the smart contract paradigm requires the community to shift focus from code-level intricacies to design-level properties. The key enabler of the shift will be a solution to the problem that we formulate as “ensuring GP’s conformance to P”, where $GP$ is the self-evolving polymorphic OO program co-developed by the community, and $P$ is the big set of properties co-defined by the same.
This paper shows that TCT is a promising pathway toward the vision. We first demonstrate it on various ERC20 tokens. By holding the transactions against intuitive design-level properties, attacks exploiting unexpected code-level anti-patterns are thwarted. We then successfully apply TCT to highly complex contract code like Uniswap-V2, exemplifying a scale of $\sim$10,000 EVM instructions per transaction across multiple contracts. The near-zero runtime overhead is a crucial contributor to TCT’s practicality. With the TCT protocol in the runtime platform, everyone can feel confident that the code-level computations conform flawlessly to the contracts’ algorithmic designs.
@@ Appendix A Appendix
This appendix provides the automatically-generated Boogie code from the “TrfrProxy” trace in Table I. It helps the understanding about Section 5 and Section 7 of the paper. The code is shown in its entirety. The “LoC of Boogie” number shown in Table I is 47, which only counts the lines after all variable declarations.
type uint256 = int;
const Zero : uint256;
axiom Zero == 0;
const TwoE8 : uint256;
axiom TwoE8 == 32768;
type address = int;
type bytes32 = int;
type uint8 = int;
const TwoE16 : uint256;
axiom TwoE16 == TwoE8 * TwoE8;
const TwoE64 : uint256;
axiom TwoE64 == TwoE16 * TwoE16 * TwoE16 * TwoE16;
const TwoE255 : uint256;
axiom TwoE255 == TwoE64 * TwoE64 * TwoE64 * TwoE16 * TwoE16 * TwoE16 * TwoE8;
const TwoE256 : uint256;
axiom TwoE256 == TwoE64 * TwoE64 * TwoE64 * TwoE64;
function evmadd(a,b:uint256) returns (uint256);
axiom (forall a,b: uint256 :: a+b < TwoE256 && a+b>=Zero ==> evmadd(a,b) == a+b);
axiom (forall a,b: uint256 :: a+b >= TwoE256 && a+b>=Zero ==> evmadd(a,b) == a+b-TwoE256);
axiom (forall a,b: uint256 :: evmadd(a,b)>=a ==> evmadd(a,b) == a+b);
function evmsub(a,b:uint256) returns (uint256);
axiom (forall a,b: uint256 :: a-b < TwoE256 && a>=b ==> evmsub(a,b) == a-b);
axiom (forall a,b: uint256 :: a-b < TwoE256 && a<b ==> evmsub(a,b) == a-b+TwoE256);
axiom (forall a,b: uint256 :: evmsub(a,b)<=a ==> evmsub(a,b) == a-b);
function evmmod(a,b:uint256) returns (uint256);
function sum(m: [address] uint256) returns (uint256);
axiom (forall m: [address] uint256, a:address, v:uint256 :: sum(m[a:=v]) == sum(m) - m[a] + v);
axiom (forall m: [address] uint256 :: ((forall a:address :: Zero<=m[a]) ==> (forall a:address :: m[a]<=sum(m))));
function nondet() returns (uint256);
var Demo.attacker1Address: [address] address;
var Demo.attacker2Address2: [address] address;
var Demo.benignUserAddress2: [address] address;
var MultiVulnToken.owner: [address] address;
var MultiVulnToken.totalSupply: [address] uint256;
var MultiVulnToken.balances: [address] [address] uint256;
var StandardToken.owner: [address] address;
var StandardToken.totalSupply: [address] uint256;
var StandardToken.balances: [address] [address] uint256;
var Token.owner: [address] address;
var Token.totalSupply: [address] uint256;
var no_reentrancy_attack._to: [address] address;
var reentrancy_attack._to: [address] address;
var reentrancy_attack.count: [address] uint256;
var _from: address;
var _to: address;
var _value: uint256;
var _fee: uint256;
procedure straightline_code ()
modifies Demo.attacker1Address, Demo.attacker2Address2, Demo.benignUserAddress2, MultiVulnToken.owner, MultiVulnToken.totalSupply, MultiVulnToken.balances, StandardToken.owner, StandardToken.totalSupply, StandardToken.balances, Token.owner, Token.totalSupply, no_reentrancy_attack._to, reentrancy_attack._to, reentrancy_attack.count, _from, _to, _value, _fee;
var tx_origin: address;
var entry_contract: address;
var BLOCKTIME: uint256;
var tmp1: uint256;
var tmp2: uint256;
var tmp3: bool;
var tmp4: bool;
var tmp5: uint256;
var tmp6: uint256;
var tmp7: uint256;
var tmp8: bool;
var tmp9: bool;
var tmp10: uint256;
var tmp11: uint256;
var tmp12: uint256;
var tmp13: bool;
var tmp14: bool;
var tmp15: uint256;
var tmp16: uint256;
var tmp17: uint256;
var tmp18: uint256;
var tmp19: uint256;
var tmp20: uint256;
var tmp21: uint256;
// def-vars
var totalSupply: uint256;
totalSupply:= MultiVulnToken.totalSupply[entry_contract];
// hypothesis
assume(Zero <= _value);
assume(_value < TwoE255);
assume(Zero <= _fee);
assume(_fee < TwoE255);
assume(totalSupply < TwoE255);
// insert invariant of entry contract
assume(forall x:address :: Zero <= MultiVulnToken.balances[entry_contract][x] && MultiVulnToken.balances[entry_contract][x] <= MultiVulnToken.totalSupply[entry_contract]);
assume(sum( MultiVulnToken.balances[entry_contract] ) == MultiVulnToken.totalSupply[entry_contract]);
tmp1:=MultiVulnToken.balances[entry_contract][_from];
tmp2:=evmadd(_fee,_value);
tmp3:= (tmp1<tmp2);
tmp4:=!tmp3;
assume(tmp4);
tmp5:=MultiVulnToken.balances[entry_contract][_to];
tmp6:=MultiVulnToken.balances[entry_contract][_to];
tmp7:=evmadd(tmp6,_value);
tmp8:= (tmp7<tmp5);
tmp9:=!tmp8;
assume(tmp9);
tmp10:=MultiVulnToken.balances[entry_contract][tx_origin];
tmp11:=MultiVulnToken.balances[entry_contract][tx_origin];
tmp12:=evmadd(tmp11,_fee);
tmp13:= (tmp12<tmp10);
tmp14:=!tmp13;
assume(tmp14);
tmp15:=MultiVulnToken.balances[entry_contract][_to];
tmp16:=evmadd(tmp15,_value);
MultiVulnToken.balances[entry_contract][_to]:=tmp16;
tmp17:=MultiVulnToken.balances[entry_contract][tx_origin];
tmp18:=evmadd(tmp17,_fee);
MultiVulnToken.balances[entry_contract][tx_origin]:=tmp18;
tmp19:=MultiVulnToken.balances[entry_contract][_from];
tmp20:=evmadd(_value,_fee);
tmp21:=evmsub(tmp19,tmp20);
MultiVulnToken.balances[entry_contract][_from]:=tmp21;
// (post) insert invariant of entry contract
assert(forall x:address :: Zero <= MultiVulnToken.balances[entry_contract][x] && MultiVulnToken.balances[entry_contract][x] <= MultiVulnToken.totalSupply[entry_contract]);
assert(sum( MultiVulnToken.balances[entry_contract] ) == MultiVulnToken.totalSupply[entry_contract]);

### v2

@@ 7.3. Complexity of the guards and math
[…]
@@
[…]
@@
Each of these guards contains the condition pair.token0 == tokenB && pair.token1 == tokenA. Due to the aforementioned duality, there are three dual-guards, not shown in Table 3. They contain the condition pair.token0 == tokenA && pair.token1 == tokenB. Hence, each pair contract needs six theorems to cover normal liquidity and swap transactions.
People may ask why the proportionality (i.e., $x_{2}/x_{1}=y_{2}/y_{1}$) and inverse-proportionality (i.e., $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$) cannot be checked concretely. There are two discrepancies between the design and the code. First, the formulas in Table 3 contain many divisions. The equality holds for real-number arithmetic, but the code performs fixed-point arithmetic. Concrete checks Figure 8 will not work due to the accumulated imprecision of fixed-point operations. Second, for every swap transaction, a small processing fee is charged by the pair as the incentive for the LPs. The rate is set by factory.swapFeeRate, which is usually set to 3 to specify a 0.3% rate. The inverse-proportionality holds only when the fee rate equals 0. Therefore, the common wisdom that “Uniswap’s swap transaction obeys the inverse-proportionality” is not precise. The actual meaning is the conjunction of two trace properties: (1) the code trace of the transaction is identical to the code trace of a zero-fee-rate transaction, (2) the latter trace, when assuming the real number arithmetic, implies the precise equality $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$. TCT successfully ensures both properties – the first is checked by the path-hash mechanism, and the second is proven by Boogie using the data type real.
7.4. Runtime overhead
The measurements in Table 4 and Table 5 are made using an HP Z4 G4 Workstation with a 3.80 GHz CPU and 32 GB RAM. Table 4 shows the runtime overhead on Geth 1.13. The Rntrcy transaction is not included because we intentionally make a false theorem for this case. We use the Golang utility time.Now().UnixNano() for the measurement. The baseline is the unmodified Geth. Every measurement is repeated using 20 transactions. Each cell reports the average and the percentage of the standard deviation.
The runtime overhead consists of the guard checking and the path hash computation. For each transaction, these two computations are repeated $10^{6}$ times for higher precision of the measurement. The overhead is the sum of these two numbers divided by the baseline. TrfrProxy and NoRntrcy are simple token transactions. The overheads are both 0.20%. As a comparison, Solythesis (li2020securing, 14) is a Solidity code instrumentation technique to insert runtime checks to enforce token properties similar to Property 1. It reports 24% slowdown in transaction processing.
Table 4. Overhead on Geth 1.13 (in nanoseconds).
@@ 7.4. Runtime overhead
Table 5. Concolic execution time and proof time.
@@
Table 5 reports the concolic execution time and the Boogie proof time. Compared to typical formal verification tasks, these numbers are quite low because the reasoning is only about a straight-line code. On the other hand, they are five orders of magnitude higher than those in Table 4. The theorem repo turns them into a one-time cost rather than a per-transaction cost. For normal transactions that the programmers anticipate, this one-time cost is paid at the testing time before the contracts process real user transactions.
7.5. Insightful violation conditions
@@
TCT’s process explicates insightful conditions at the detailed code level, such as “tx.origin != pair” and “to != pair” for Swap. The former states that the swap transaction must not be issued by the pair itself, and the latter states that the account receiving the swapped-out token from the pair must not be the pair itself. When the conditions are violated, the inverse proportionality will not hold. To show this, we issue six consecutive swap transactions to Geth. The reserves of the pair (i.e., $x$ and $y$) are shown in Table 6. The first three transactions conform to to != pair, and the last three violate it. The deviation is the current $x\times y$ divided by the previous $x\times y$ then subtracting 1. A perfect inverse proportionality implies zero deviation. Because of the 0.3% fee, the deviations of the first three transactions are [0.0008, 0.0010], which are reasonable. However, for the last three transaction, $y$ becomes unchanged because the amount of the swapped-out token goes back to the pair. This makes the deviation very large (in fact, arbitrarily large). Hence, if an investment strategy depends on Uniswap unconditionally ensuring inverse proportionality, it can be dangerous.
Table 6. Inverse proportionality violated when to=pair.
condition $x$ $y$ $x\times y$ Deviation
to $\neq$ pair 2000 1000 2000000 N/A
to $\neq$ pair 2010 996 2001960 0.0010
to $\neq$ pair 2030 987 2003610 0.0008
to = pair 2050 987 2023350 0.0099
to = pair 2070 987 2043090 0.0098
to = pair 2190 987 2161530 0.0580
8. Roadmap Ahead
Our prototype shows that TCT points to a promising direction for faithful and scalable runtime verification, but there are future work items to be discussed when we consider the real-world deployment.
8.1. Consensus of theorem-proving
@@
The Ethereum nodes need to have a consensus about which theorems have been proven. Boogie’s underlying SMT prover is Z3 (z3, 40), which is deterministic (unless timeout-based proving tactics are enabled). Z3 gives three possible results: “sat”, “unsat” and “timeout”. Determinism implies that a group of honest nodes will not have a “sat vs. unsat” split, although some nodes may time out. Note that some slow nodes falling behind the pace of the majority is a normal situation in Ethereum’s operation, which does not prevent the consensus to be formed. Alternatively, we can design an explicit prover-gas mechanism, similar to the existing gas mechanism, to calculate the amount of Z3 computation. It ensures that all honest nodes will yield an identical result: “sat”, “unsat” or “out of prover gas”. Yet another mechanism to consider is to create a dedicated prover-node consortium, which forms the consensus about which theorems have been proven. Normal Ethereum nodes no longer need to run Z3. Their theorem repositories only need to sync with the latest published snapshot of the consortium’s repository.
8.2. Specifying and reasoning about summaries
Reference (Compositional-testing, 8) introduces the notion of summaries for symbolic execution. Basically, a trace does not need to get into a piece of code (e.g., a branch, a loop or a function) if a postcondition of the code is sufficient to prove a property. For example, the branch “if (b) a=1; else a=2;” can be summarized as “modifies(a); assume(a==1 || a==2);”, meaning that no variable other than a is modified, and a is set to either 1 or 2. Similarly, a loop that calculates the sum of an array or initialize an array’s elements can be summarized so that the loop can be skipped. We believe that summaries are particularly useful for basic library functions.
8.3. Carrying multiple theorems
Currently, every transaction only carries one theorem. It can be extended to multiple theorems for a transaction shown in Figure 9. Consider a batch transaction in which ContractA::BatchFoo has a loop to call Foo in callee contracts (e.g., tokens). The transaction can carry theorem2, theorem3 and theorem4 to prove the postconditions of the three functions. Based on our case-studies, it is not very obvious what safety properties are meaningful to define for the entire batch (because the simple act of batching does not seem to introduce a safety concern). However, in case a postcondition is defined for ContractA::BatchFoo, the symbolic execution can use the postconditions of the callee functions as their summaries to skip the function bodies to prove theorem1. Essentially, the entire trace is treated as four traces to avoid the combinatorial complexity.
Figure 9. A transaction to be proven by multiple theorems.
9. Related Work
PCC and PCA. Proof-Carrying Code (PCC) (necula1996safe, 18) is a technology for a code module from an untrusted producer (e.g. a device driver) to carry its own safety proof, so that a trusted party (e.g.OS kernel) can accept the module with confidence. Proof-Carrying Authentication (PCA) (PCA, 2) is an authentication mechanism inspired by PCC. Real-world authentication policies, such as those involving “certificate authority”, “delegate”, the “speaks-for” relation, etc., are essentially statements about statements, so the authentication logic is higher-order, which is undecidable. PCA requires every request to carry a proof of its legitimacy based on the policies, so the server only needs a decidable proof-checker. PCA emphasizes its logic extensibility. Anyone can define operations and prove lemmas about them. Once the proofs are confirmed, the server accepts the lemmas into the logic for future proofs. This can be analogous to TCT’s theorem repository mechanism. Between PCC and PCA, the TCT concept is closer to the latter – a theorem attests to the safety of every runtime request rather than a code module.
@@ 9. Related Work
Static verification for smart contracts. Oyente (luu2016making, 16) and Mythril (mueller2018smashing, 17) are symbolic verifiers for EVM bytecode. Solidity SMT Checker (Solidity-SMTChecker, 32) is a feature built into the Solidity compiler to check assertions and alert possibilities of reentrant calls. Halmos (park:halmos, 22) is a recent symbolic analyzer for EVM bytecode compiled from Solidity. In our prototype, we consulted the Halmos code for the EVM semantics. VerX (permenev2020verx, 24) is a static verifier to prove temporal properties for a set of contracts, with certain assumptions about the “external” contracts that they interact with. The temporal logic is focused on properties that correlate contract states across different transactions. It can reason about liveness properties, which TCT is unable to. Annotary (weiss2019annotary, 38) is a Mythril-based verifier that can examine cross-contract transactions. To tackle the state explosion problem, Annotary uses different strategies to treat different variables as either concrete or symbolic, depending on their data types. The authors refer to this approach as “concolic execution”, very different from ours.
Static verification considers one or a few “contracts of interest” as the program, and makes assumptions about all other contracts and hypothetical malicious users. TCT is a runtime verification that treats all contracts as a single GP without these assumptions.
Runtime verification for smart contracts. OpenZeppelin offers the SafeMath library (safe-math, 19) to throw runtime exceptions upon integer overflows. The Solidity compiler v0.8.0+ generates these assertions in the EVM code by default, but also introduces the unchecked statement to selectively omit these assertions. We observe that the unchecked statement is used fairly often, perhaps mainly for gas-saving. However, there are scenarios where an overflow is intentional. About the reentrancy bugs, Grossman et al. observe that “callbacks are not esoteric” in real-world contracts (ECF_reentrancy, 10), so a precise notion should be defined to tell which callbacks are “bugs”. They define a notion Effectively Callback Free (ECF), analogous to serializability of concurrent systems. A commutativity-check mechanism is built into EVM, which determines whether a nested-call trace intrinsically needs to be nested, or it is serializable because the nested calls are logically commutative. Rodler et al. develop another mechanism called Sereum (Sereum_reentrancy, 27). It taint-tracks storage variables that affect the control flow, and locks them to prohibit further updates when the victim contract is entered again deeper in the call stack. The motivation of TCT is complementary to these techniques. We target interface specifications, intentionally agnostic to low-level bug patterns. On the other hand, the assertions in SafeMath, the serializability-like ECF checking and the taint-based Sereum checking are all about trace properties, which should be verifiable symbolically. In this sense, TCT does not exclude the ability to verify low-level properties.
@@
The Solythesis approach (li2020securing, 14) is more related to our goal toward interface specifications. It targets aggregation properties, such as sum-of-map, count-of-subset, etc. It instruments Solidity code to insert incremental updates so that aggregations can be calculated with much lower cost than naive computations. The instrumented code adds auxiliary arrays and maps to keep track of the incremental updates. It is unclear whether Solythesis handles multi-contract situations, as the evaluation is only conducted on 23 individual token contracts. The instrumented contracts show a 24% slowdown, two orders of magnitude higher than that of TCT.
10. Conclusion
The uniqueness of our work is to view the entire set of smart contracts as a polymorphic gigantic program $GP$, and make the $GP$ the subject of verification. TCT combines the benefits of runtime checking and symbolic verification – it is faithful to the $GP$ like runtime checking, but enables interface specifications to be symbolically verified. Because of its concolic nature, programmers only need to explicate guard conditions according to their intentions, while leaving attackers obligated to prove whether unexpected conditions and code paths also maintain the interface specifications. This greatly reduces the proof obligation, making TCT significantly scalable. The Uniswap scenario exemplifies a scale of $\sim$10,000 EVM instructions per transaction. Faithful symbolic verification for smart contract code at this complexity level has not been achieved before. The near-zero runtime overhead is crucial for TCT’s practicality.
TCT’s goal is complementary to compile-time static verification. With the faithfulness, interface specifications are an inseparable part “engraved” in the source code, i.e. people can treat them as contractual promises in the same manner as the rest of the source code. We hope this strong assurance will incentivize the Ethereum community to shift the focus from detecting code-level intricacies to defining interface specifications.
@@ Appendix A Appendix
A.1. Axioms of mathematical operations
LABEL:code:axiommath lists some axioms in TCT’s basic library. They are not specific to the discussed case. In the axioms, m is a map from address to int. Notation m[a:=v] denotes a map that is almost identical to m, but its index a contains the value v.
function sum(m: [address] int) returns (int);
axiom forall m: [address] int, a:address,v:int:: sum(m[a:=v]) == sum(m)-m[a]+v;
axiom forall m: [address] int :: ((forall a:address :: 0 <= m[a]) ==> (forall a:address :: m[a] <= sum(m)));
function evmadd(a,b:int) returns (int);
axiom forall a,b:uint256 :: a+b < 2^256 && a+b >= 0 ==> evmadd(a,b) == a+b;
axiom forall a,b: uint256 :: a+b >= 2^256 ==> evmadd(a,b) == a+b-2^256;
function evmsub(a,b:int) returns (int);
axiom forall a,b: uint256 :: a-b < 2^256 && a >= b ==> evmsub(a,b) == a-b;
axiom forall a,b: uint256 :: a < b ==> evmsub(a,b) == a-b+2^256;
Code 8: Axioms of mathematical operations.
A.2. The protocol flow of the repo miss scenario
Similar to the protocol flows in Figure 4, the flow corresponding to the theorem-repo miss scenario is shown in Figure 10.
Figure 10. The protocol flow of the repo miss scenario.
A.3. Additional details about concolic execution and SVT
In Section 5.2, we show an EVM code segment with its initial stack and final stack, which is shown again in Figure 11. Some clarifications are helpful in understanding the EVM code and the SVT nodes.
  • The EVM storage is a key-value pair. Both the key and the value are 32-byte. The key is obtained by computing a SHA3 over the slot number of the map variable and the index of the element according to reference (soliditylayout, 29). In our example, the slot number is 2 (i.e., 878 PUSH1 02), which corresponds to the map variable balances based on the compilation symbol info of MultiVulnToken. The index is Partial32B((12,31), _from). Hence, the MapElement subtree is equivalent to balances[_from] at the Solidity level.
Figure 11. Execution of an EVM code trace segment.